OpenStudy (anonymous):

Notice a pattern: 3^2 + 4^2 = 5^2 5^2 + 12^2 = 13^2 7^2 + 24^2 = 25^2 9^2 + 40^2 = 41^2 Conjecture a pattern, test it. Prove it

OpenStudy (anonymous):

3:4:5 is the ratio of a perfect triangle it holds true for the rest of the numbers too.

OpenStudy (anonymous):

its like : (1+4)^2, then (5+8)^2 and then (13+12)^2 and (25+16)^2. can you see the patern now?

OpenStudy (anonymous):

this is what i see: \[(2n + 1)^2 + \left( 4\sum_{i = 1}^{n}i \right)^2 = \left( 1 + 4\sum_{i = 1}^{n}i \right)^2\] appears to work. which class is this for btw? might narrow down how we're supposed to solve it. give feedback plz ^_^